# Probabilistic Decomposed Linear Dynamical Systems for Robust Discovery of Latent Neural Dynamics

### v1

@@ Abstract
Time-varying linear state-space models are powerful tools for obtaining mathematically interpretable representations of neural signals. For example, switching and decomposed models describe complex systems using latent variables that evolve according to simple locally linear dynamics. However, existing methods for latent variable estimation are not robust to dynamical noise and system nonlinearity due to noise-sensitive inference procedures and limited model formulations. This can lead to inconsistent results on signals with similar dynamics, limiting the model’s ability to provide scientific insight. In this work, we address these limitations and propose a probabilistic approach to latent variable estimation in decomposed models that improves robustness against dynamical noise. Additionally, we introduce an extended latent dynamics model to improve robustness against system nonlinearities. We evaluate our approach on several synthetic dynamical systems, including an empirically-derived brain-computer interface experiment, and demonstrate more accurate latent variable inference in nonlinear systems with diverse noise conditions. Furthermore, we apply our method to a real-world clinical neurophysiology dataset, illustrating the ability to identify interpretable and coherent structure where previous models cannot.
@@ 2 Background and Related Work
$\bm{y}_{t}$ $=\bm{D}\bm{x}_{t}+\bm{\epsilon}_{y_{t}},$ $\quad\bm{\epsilon}_{y_{t}}\sim\mathcal{N}(0,\bm{\Sigma}_{y}),$ $\quad{\rm(Observations)}$ (1)
@@
where $\bm{x}_{t}\in\mathbb{R}^{N}$ is the latent state, $f_{t}(\cdot)$ is the dynamics function, and $\bm{D}\in\mathbb{R}^{M\times N}$ is a linear observation function. Our work focuses on the case when $N<M$, which compresses high-dimensional signals into a low-dimensional latent space. By choosing $f_{t}$ to be a time-varying linear operator, we can approximate complex nonlinear dynamics with simple locally-linear components, balancing expressivity with mathematical interpretability. However, learning a time-varying linear operator from data can be challenging, and typically requires additional constraints on the underlying generative model to identify meaningful representations.
@@ 3.2 Probabilistic Time-Informed Sparsity
In decomposed models, we aim to achieve two goals simultaneously: sparsity and smoothness of coefficients over time. Motivated by this, we incorporate dynamics-informed probabilistic structure. First, we assume that each coefficient evolves independently of the others. Second, we introduce a hierarchical variance parameter $\gamma_{t,k}$ that controls the sparsity for each $c_{t,k}$. Large values of $\gamma_{t,k}$ encourage coefficients to be active while small values effectively shrink them to zero. Moreover, we introduce dynamics information during sparse inference by encouraging a similar active support set in consecutive time slices through the variance hyperpriors. Put together, the resulting coefficient transition density in p-dLDS becomes,
@@
The density $p(c_{t,k}|c_{t-1,k}\gamma_{t,k})$ is factorized as a product of two independent priors, $\mathcal{N}(c_{t-1,k},\sigma_{t-1,k}^{2})\mathcal{N}(0,\gamma_{t,k})$, where the first prior encourages dynamic smoothness while the second encourages sparsity. The second density in equation (5) is defined similarly to the hyperprior in SBL-DF (i.e., $p(\gamma_{t,k}|c_{t-1,k})=\mathcal{IG}(\xi,\xi c_{t-1,k}^{2})$ where $\xi$ weighs the influence of the dynamics when estimating $\gamma_{t,k}$). The resulting graphical model is shown in Figure  1B.
@@ 3.3 Inference and Learning
$p(\bm{x},\bm{y},\bm{c},\bm{\gamma}|\theta)=p(\bm{x}_{1})\left[\prod_{t=1}^{T}p(y_{t}|x_{t})\right]\left[\prod_{t=1}^{T-1}p(\bm{x}_{t+1}|\bm{x}_{t},\bm{c}_{t})\left[\prod_{k=1}^{K}p(c_{t+1,k}|c_{t,k},\gamma_{t+1,k})p(\gamma_{t+1,k}|c_{t,k})\right]\right],$ (6)
@@
Given our decomposition equation (4), the optimal coordinate ascent approximate posterior becomes
$q(\bm{x})=\mathcal{N}(\bm{x}_{1};\bm{\mu}_{1},\Sigma_{1})\left[\prod_{t=2}^{T}\mathcal{N}(\bm{y}_{t};\bm{D}\bm{x}_{t},\bm{\Sigma}_{y})\right]\left[\prod_{t=1}^{T}\mathcal{N}(\bm{x}_{t};\bm{l}_{t-1}+\bm{F}_{t-1}\bm{l}_{t-1}+\bm{b}_{t},\bm{\Sigma}_{x})\right],$ (8)
@@ 4.1 Synthetic Dynamical Systems
NASCAR with Random Speeds. We evaluate our inference procedure on the NASCAR dataset [22, 28, 20]. Since this system is easily mean-centered, we can isolate the effect of our proposed inference procedure as offset terms are not necessary. To make this dataset more realistic, we introduce speed variability into the ground truth dynamics as opposed to having a perfect constant speed at all time. Specifically, whenever a trajectory enters a new segment of the track, the system experiences a random change in speed. We trained all models on 30 trials, each consisting of 1000 time steps (Fig. 2A) with randomly sampled initial points, and a randomly constructed 10-dimensional linear observation matrix (see Appendix C.1 for full details). Performance is evaluated on 30 held-out trials where the ground truth switching states are defined by the different segments of the track. In all models, we set the $M=10$, $N=2$ and $K=4$ for DOs or switching states.
@@ Appendix A Appendix / supplemental material
     $\bm{F}_{t}\leftarrow\sum_{k=1}^{K}\bm{f}_{k}\widehat{c}_{k,t}$
     $\bm{l}_{1:T},\bm{\Sigma}_{x}\leftarrow{\rm KalmanSmoother}(\bm{\widehat{l}}_{1:T},\bm{F}_{1:T},\theta)$
@@ F.1 Hyperparameter Settings
For p-dLDS, the relevant hyperparameters consists of the SBL-DF dynamics tradeoff $\xi$, and the offset window size $S$. In the NASCAR experiment, we isolate the effect of the probabilistic inference procedure by setting $S=T$, removing the influence of the time-varying offset term. For $\xi$, we perform a random search by uniformly sample over the log of the interval $[10^{-3},10^{3}]$ and found that $\xi=0.945$ was optimal. For the Lorenz experiment, we also optimize for the window size $S$ by uniformly sample a discrete index on the interval $\{2,\dots,T\}$. For the Lorenz experiment, the optimized hyperparameters are $S=85$ and $\xi=8.928$.

### v2

@@ Abstract
Time-varying linear state-space models are powerful tools for obtaining mathematically interpretable representations of neural signals. For example, switching and decomposed models describe complex systems using latent variables that evolve according to simple locally linear dynamics. However, existing methods for latent variable estimation are not robust to dynamical noise and system nonlinearity due to noise-sensitive inference procedures and limited model formulations. This can lead to inconsistent results on signals with similar dynamics, limiting the model’s ability to provide scientific insight. In this work, we address these limitations and propose a probabilistic approach to latent variable estimation in decomposed models that improves robustness against dynamical noise. Additionally, we introduce an extended latent dynamics model to improve robustness against system nonlinearities. We evaluate our approach on several synthetic dynamical systems, including an empirically-derived brain-computer interface experiment, and demonstrate more accurate latent variable inference in nonlinear systems with diverse noise conditions. Furthermore, we apply our method to a real-world clinical neurophysiology dataset, illustrating the ability to identify interpretable and coherent structure where previous models cannot. 111 Code is available at:
https://github.com/siplab-gt/probabilistic-decomposed-linear-dynamical-systems
@@ 2 Background and Related Work
$\bm{y}_{t}$ $=\bm{D}\bm{x}_{t}+\bm{d}+\bm{\epsilon}_{y_{t}},$ $\quad\bm{\epsilon}_{y_{t}}\sim\mathcal{N}(0,\bm{\Sigma}_{y}),$ $\quad{\rm(Observations)}$ (1)
@@
where $\bm{x}_{t}\in\mathbb{R}^{N}$ is the latent state, $f_{t}(\cdot)$ is the dynamics function, and $\bm{D}\in\mathbb{R}^{M\times N}$ and $\bm{d}\in\mathbb{R}^{M}$ describe a linear observation function. Our work focuses on the case when $N<M$, which compresses high-dimensional signals into a low-dimensional latent space. By choosing $f_{t}$ to be a time-varying linear operator, we can approximate complex nonlinear dynamics with simple locally-linear components, balancing expressivity with mathematical interpretability. However, learning a time-varying linear operator from data can be challenging, and typically requires additional constraints on the underlying generative model to identify meaningful representations.
@@ 3.2 Probabilistic Time-Informed Sparsity
In decomposed models, we aim to achieve two goals simultaneously: sparsity and smoothness of coefficients over time. Motivated by this, we incorporate dynamics-informed probabilistic structure. First, we assume that each coefficient evolves independently of the others. Second, we introduce a hierarchical variance parameter $\gamma_{t,k}$ that controls the sparsity for each $c_{t,k}$. Moreover, we introduce dynamics information during sparse inference by encouraging a similar active support set in consecutive time slices through the variance hyperpriors. Put together, the resulting coefficient transition density in p-dLDS becomes,
@@
We define the first term on the right-hand side with the following functional form,
$p(c_{t,k}|c_{t-1,k},\gamma_{t})\propto\exp\left(-\frac{c_{t,k}^{2}}{2\gamma_{t,k}}-\frac{(c_{t,k}-c_{t-1,k})^{2}}{2\sigma^{2}_{t-1,k}}\right)\propto\mathcal{N}(c_{t-1,k},\sigma_{t-1,k}^{2})\mathcal{N}(0,\gamma_{t,k}).$ (6)
This density captures the constraints of sparsity and smoothness for the inferred coefficients $c_{t,k}$. When the variance around zero $\gamma_{t,k}$ is small, this structure promotes sparsity by shrinking coefficient values towards zero. Conversely, when the variance around the previous time step $\sigma^{2}_{t-1,k}$ is small, it encourages smooothness by shrinking coefficients towards the previous value. While the idea of combining two shrinkage effects in a single density has been explored in previous works [15, 5, 24, 18], those approaches generally require manual balancing of the two penalties. In contrast, we devise a procedure in the following section that estimates these variance parameters automatically during inference and learning.
The second density on the right-hand side from equation (5) is defined similarly to the hyperprior in SBL-DF. (i.e., $p(\gamma_{t,k}|c_{t-1,k})=\mathcal{IG}(\xi,\xi c_{t-1,k}^{2})$ where $\xi$ weighs the influence of the dynamics when estimating $\gamma_{t,k}$). The resulting graphical model is shown in Figure  1B. We note that since the value of the previous coefficient is squared, the overall prior placed on the inverse gamma density follows a $\chi^{2}$ distribution.
@@ 3.3 Inference and Learning
$\begin{split}p(\bm{x},\bm{y},\bm{c},\bm{\gamma}|\theta)=p(\bm{x}_{1})&\left[\prod_{t=1}^{T}p(\bm{y}_{t}|\bm{x}_{t})\right]\\ &\left[\prod_{t=1}^{T-1}p(\bm{x}_{t+1}|\bm{x}_{t},\bm{c}_{t})\left[\prod_{k=1}^{K}p(c_{t+1,k}|c_{t,k},\gamma_{t+1,k})p(\gamma_{t+1,k}|c_{t,k})\right]\right],\end{split}$ (7)
@@
Our assumed decomposition in equation (4) allows us to define the latent state transition density as $p(\bm{x}_{t+1}|\bm{x}_{t},\bm{c}_{t})=p(\bm{x}_{t+1}=\bm{l}_{t+1}+\bm{b}_{t+1})=\mathcal{N}(\bm{x}_{t+1};\bm{l}_{t}+\bm{F}_{t}\bm{l}_{t}+\bm{b}_{t})$. Substituting this into equations (7) and (3.3), we get that the optimal coordinate ascent approximate posterior becomes,
$q(\bm{x})=\mathcal{N}(\bm{x}_{1};\bm{\mu}_{1},\Sigma_{1})\left[\prod_{t=2}^{T}\mathcal{N}(\bm{y}_{t};\bm{D}\bm{x}_{t}+\bm{d},\bm{\Sigma}_{y})\right]\left[\prod_{t=1}^{T}\mathcal{N}(\bm{x}_{t};\bm{l}_{t-1}+\bm{F}_{t-1}\bm{l}_{t-1}+\bm{b}_{t},\bm{\Sigma}_{x})\right],$ (9)
@@ 4.1 Synthetic Dynamical Systems
NASCAR with Random Speeds. We evaluate our inference procedure on the NASCAR dataset [26, 32, 23]. Since this system is easily mean-centered, we can isolate the effect of our proposed inference procedure as offset terms are not necessary. To make this dataset more realistic, we introduce speed variability into the ground truth dynamics as opposed to having a perfect constant speed at all time. Specifically, whenever a trajectory enters a new segment of the track, the system experiences a random change in speed. We trained all models on 30 trials, each consisting of 1000 time steps (Fig. 2A) with randomly sampled initial points, and a randomly constructed 10-dimensional linear observation matrix (see Appendix C.1 for full details). Performance is evaluated on 30 held-out trials where the ground truth switching states are defined by the different segments of the track. In all models, we set the $M=10$, $N=2$ and $K=4$ for DOs or switching states. In our experiments, we define the "discrete states" for decomposed models as the DO state with the largest coefficient magnitude.
@@ 5 Conclusion
Acknowledgements and Disclosure of Funding Sources
Y.C. and C.R. were funded by the James S. McDonnell Foundation (grant number 22002039), with Y.C. being further funded by National Institutes of Health (grant number 2T32EB025816), and C.R. being further funded by the Julian T. Hightower Chair. Y.C. and K.J. were part of the Georgia Tech/Emory NIH/NIBIB Training Program in Computational Neural Engineering (T32EB025816). N.M. was funded by The Kavli Foundation NeuroData Discovery award. A.S.C. were partially supported by the NSF CAREER Award (2340338) and a Johns Hopkins Bridge Grant. S.A. is supported by the National Center for Advancing Translational Sciences of the National Institutes of Health (Award Number UL1TR002378 and KL2TR002381). The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
@@ Appendix A Appendix / supplemental material
$\widehat{\bm{x}}_{t}\leftarrow{\bm{D}}^{+}\bm{y}_{t}$triangleright Initialize latent state with PCA
@@
     $\bm{\widehat{F}}_{t}\leftarrow\sum_{k=1}^{K}\bm{f}_{k}\widehat{c}_{k,t}$
     $\bm{l}_{1:T},\bm{\Sigma}_{x}\leftarrow{\rm KalmanSmoother}(\bm{y}_{1:T},\bm{\widehat{b}}_{1:T},\bm{\widehat{F}}_{1:T},\theta)$
@@ F.1 Hyperparameter Settings
For p-dLDS, the relevant hyperparameters consists of the SBL-DF dynamics tradeoff $\xi$, and the offset window size $S$. We use random search with a budget of 1000 samples to determine the values of $S$ and $\xi$ and fit a separate model for each set of hyperparameters. In the NASCAR experiment, we isolate the effect of the probabilistic inference procedure by setting $S=T$, removing the influence of the time-varying offset term. For $\xi$, we perform a random search by uniformly sample over the log of the interval $[10^{-3},10^{3}]$ and found that $\xi=0.945$ was optimal. For the Lorenz experiment, we also optimize for the window size $S$ by uniformly sample a discrete index on the interval $\{2,\dots,T\}$. For the Lorenz experiment, the optimized hyperparameters are $S=85$ and $\xi=8.928$. For the real dataset, the optimal offset is $S=76$ which is smaller than the timescale of p-dLDS coefficient switching (around 150 time points), suggesting that the same DO dynamics may persist even as the fixed points of the system fluctuates throughout the experiment.
@@ Appendix H Limitations
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We propose a probabilistic treatment and an extended dynamics formulation in decomposed models (Section 3). We demonstrate that these changes reduce estimation errors and finds coherent structure where previous models fail in many challenging synthetic examples, and a noisy real-world example (Section 4).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: A limitations section is provided in Appendix H due to space constraints.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide proofs of our lemmas in B.1 and B.2.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide synthetic data details in Appendix C, metric definitions in Appendix D, and hyperparameter details in Appendix F.1. Moreover, we release our code with our submission.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide functions to generate the synthetic datasets from our experiments. Unfortunately, we are unable to release the neurophysiological dataset due to request from our collaborators.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We specify that all datasets are split 50:50 for train and test in the main paper. Moreover, we provide optimizer and hyperparameter settings in Appendix F.1.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We provide standard deviations in Appendix E.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide details about hardware in Appendix F.2.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Most of our datasets are synthetically generated or derived from publicly available sources, which we refer to throughout the paper. For our real-world experiment, we include the IRB identifier.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our proposed model aims to reveal coherent patterns in time-series data, serving as a scientific tool similar to PCA. As a probabilistic technique not designed for content generation or automated decision-making, we do not anticipate direct societal impacts.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our method does not generate synthetic content that is at risk of abuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification:
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We include details on where we obtained our data (either synthetically generated, or part of an existing dataset) in the main text and appendix. We also point to the commonly package used for SLDS and rSLDS in the appendix.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: We provide the IRB identifier in Appendix G which includes the informed consent form with experiment details.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: We include the IRB identifier number in Appendix G, but removed any information related to the institution conducting the experiments.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.